i can not solve this inequality please sir how to solve this inequality
Ix+1I+Ix-2I

Grade:12th pass

5 Answers

DR STRANGE
7 years ago
Ix+1I+Ix-2I = 2x-1 ….....for x>2
                    =x+1 -(x-2) = 3 …......for -1
                     = -(x+1) -(x-2) = 1-2x ….......for x
DR STRANGE
7 years ago
mistake...............................................................
=3 for...... -1
= 1-2x ….for x
DR STRANGE
7 years ago
i am trying to write conditions in the answer but after posting the conditions are vanishing , you can see above, dont know why is it happening.... :(
somi teez
6 years ago
hello,
I x+1 I +Ix-2I
there are two break points -1 and 2
case 1: x
case 2: -1
case 3:x>2              modulus opens with positive (x+1)+(x-2)=2x-1
hence we have these values at all possible cases
 
Hari Sundar Pyda
6 years ago
The answer is shown below-Ix+1I+Ix-2IWhen x2,x+1+x-2= 2x-1When 1